OpenStudy (lovelycharm):

3x + 2y = 17 2x + 3y = 26 elimination

OpenStudy (anonymous):

multiply the first equation by 2 and the second equation by -3. This will cancel out the x's and then you can solve for y. Can you do that ?

OpenStudy (anonymous):

to find the value of y,multiple the 1st equation by -2 and 3 to the 2nd equation.cancel out the x then add the values in the y and the constant.to find the value of x multiply the 1st equation by negative 3 and the 2nd equation by 2 then cancel out the y and add the remaining x and the constant.

OpenStudy (lovelycharm):

i dont get it go u write it down for me @jedai17

OpenStudy (anonymous):

3x + 2y = 17 --> (2) 3x + 2y = 17 2x + 3y = 26 --> (-3)2x + 3y = 26 ----------------- 6x + 4y = 34 (result of multiplying by 2) -6x - 9y = - 78 (result of multiplying by -3) -----------------add - 5y = - 44 -5y/-5 = - 44/-5 y = 44/5 = 8 4/5 = 8.8 now sub 8.8 in for y in either of the original equations.. 3x + 2y = 17 3x + 2(8.8) = 17 3x + 17.6 = 17 3x = 17 - 17.6 3x = - .6 3x/3 = - .6/3 x = -.2 check.... 2x + 3y = 26 2(-.2) + 3(8.8) = 26 -.4 + 26.4 = 26 26 = 26 (correct) answer : x = - .2 and y = 8.8 Do you understand ?

OpenStudy (lovelycharm):

thank u @kelliegirl33

OpenStudy (anonymous):

it not a problem.....I was glad to help :)